# The Streaming Batch Model for Efficient and Fault-Tolerant Heterogeneous Execution

### v4

@@ 5.2.1 ResNet Training
Compared to single-node ML data loaders, Ray Data offers: (1) failure isolation between heterogeneous resources, and (2) ability to leverage heterogeneous clusters.
5.2.2 Pre-Training Stable Diffusion
Pre-training of large ML models is one of the most demanding heterogeneous workloads. We run the Stable Diffusion (SD) pre-training pipeline shown in Figure 1(b) and compare different execution modes. We execute 1 training epoch over a dataset of 2 billion images, on a cluster of 4×4$\times$4 × p4de.4xlarge nodes, with 8 A100 GPUs on each node. This pipeline is challenging because it requires both CPUs and GPUs for data preprocessing: (1) loadText/loadImage+clip (CPU): Load pairs of image and text, perform preprocessing, (2) Encoder (GPU): Use a pre-trained encoder model, one for images and one for text, to produce dense embeddings, and (3) UNet.train() (GPU): Train SD on the embeddings.
Figure 8 shows training throughput and total cost. PyTorch DL is a data loader custom-built for PyTorch that statically partitions work on a process pool. Its throughput is lowest because Encoder preprocessing competes with trainers for GPU memory. Ray Data-staged emulates batch processing by running data preprocessing as an offline job and storing precomputed embeddings in cloud storage. This is preferable if the same embeddings are used multiple times. Ray Data-staged achieves 19% higher throughput because UNet is given full GPUs.
Ray Data runs all data preprocessing concurrently with training. This is preferable if using random transforms or for iterative development. Ray Data also leverages heterogeneous clusters, placing Encoders on smaller A10G GPUs (g5.2xlarge). This results in 31% better throughput than PyTorch DL, because UNet has full GPU resources, and 15% better throughput than Ray Data-staged, because embeddings are kept in memory.
Takeaways:
Compared to existing ML data loaders, Ray Data can: (1) be used for both batch and online data preprocessing, and (2) leverage heterogeneous distributed execution.
Figure 9: Synthetic benchmark run times for systems under different memory limits. Grey means the system is unable to finish due to OOM. Ray Data(-Part.) means Ray Data without streaming repartition. Ray Data(-Adapt.) means Ray Data without adaptive memory-aware scheduling.
Figure 10: Effect of partition sizes on throughput in Ray Data.
5.3 Microbenchmarks
5.3.1 Memory-aware scheduling
We evaluate how batch, stream, and streaming batch systems schedule heterogeneous pipelines with and without memory pressure. The 3-stage pipeline is: (1) Load (CPU): 160 tasks, each producing 500 1 MB rows after 5s, (2) Transform (CPU): sleep for 0.5s per row, then return a different 1 MB row, (3) Inference (GPU): 0.5s per batch of 100 rows. We use 1 m6i.2xlarge node with 8 vCPUs, 4 simulated GPU slots, and 32 GB RAM. The theoretical best job completion time with unlimited memory is $(160\times 5s+800\times 0.5s)/8=150s$.
Figure 10 shows job completion time vs. total memory limit. We limit memory through system-specific configurations, e.g., executor memory for Spark. We also use POSIX rlimit to verify that each system respects its memory limit, and tune each system’s parallelism (e.g., executor count) if not.
Spark materializes all data between stages, achieving at best 2.35×2.35$\times$2.35 × optimal run time. At 12–14GB memory, Spark must use fewer executors resulting in 4.34×4.34$\times$4.34 × the optimal run time, and at lower memory limits, Spark is unable to finish. This is because Spark requires static partitioning (§ 2.2), and the initial number of Load tasks produces too-large partitions.
Flink is less sensitive to the memory limit than Spark and achieves up to 1.68×1.68$\times$1.68 × optimal. This is because executors dynamically materialize output partitions to avoid running out of memory (§ 2.3). At lower memory limits, Flink must run fewer executors because it uses multithreading and slot sharing to multiplex a CPU slot among physical operators, making executors vulnerable to OOM under memory pressure. This results in up to 2×2$\times$2 × worse throughput.
We also compare against tf.data because unlike PyTorch DL, it offers an adaptive scheduler and memory budget, similar to Ray Data. However, we found that the memory budget was not always enforced, requiring manual tuning of the thread count. tf.data achieves the same throughput as Ray Data at 16 GB memory limit, but is unable to finish at lower memory limits.
Ray Data is able to finish in 1.3×1.3$\times$1.3 × the optimal run time at all memory limits except the lowest, due to Ray Data’s streaming repartition (Figure 3(b)) and adaptive scheduler (§ 4.3). We further conduct ablation studies on Ray Data’s performance. Ray Data(-Part.) disables Ray Data’s streaming repartition, resulting in too-large initial partitions similar to Spark. Ray Data(-Adapt.) disables Ray Data’s adaptive scheduler, resulting in a conservative policy similar to Figure 4(a) and 10–88% worse performance than Ray Data. Ray Data is also less sensitive than Flink to memory pressure because the system explicitly time-slices executors at task granularity, instead of using multithreading.
Takeaways:
For heterogeneous applications under memory pressure, batch processing systems are unstable. Ray Data is as stable as Flink, due to its streaming repartition, and also more adaptive, thanks to its scheduler.
@@ 5.3.2 Overhead of partitioning
Compared to stream processing, Ray Data also uses dynamically sized partitions, but with a centralized scheduler. A possible concern is the system overhead per partition. We evaluate the impact of partition number on throughput in Ray Data with a 2-stage synthetic pipeline. We use 8192×8192$\times$8192 ×1 MB input rows, and simulate 10 ms processing time per row per stage. Figure 10 shows the throughput vs. partition size. The smallest partition sizes incur overhead from RPCs and bookkeeping, and the largest result in poor load-balancing. To strike a balance, Ray Data’s default target partition size is 128 MB.
@@ 6 Related Work
Recent efforts to unify the batch and stream processing models include Apache Beam [4] and Google Cloud Dataflow [3]. They focus on providing a unified API layer, rather than unifying the execution model. Attempts at execution model unification include Naiad [28], Flink’s batch execution mode [9], and Spark Streaming [47]. Naiad shows that the stream processing model is suitable for producing results both incrementally and in bulk, but it does not support dynamic parallelism reconfiguration. Like other batch processing systems, Flink’s batch execution mode executes one stage at a time.
Spark Streaming partitions the input stream into discretized “microbatches”, each executed as a distinct Spark job. Drizzle [43] improves on Spark Streaming’s latency without sacrificing adaptability, but suffers the same flaws for heterogeneous pipelines: pipelining across heterogeneous resources would require a custom inter-job Spark scheduler, and the data partitioning within a microbatch is static. Also, since Spark requires stateless tasks, it imposes high overheads in ML inference and training from repeatedly loading the model into GPU memory for each task. Petastorm [17] is a Python library that bridges Spark’s data processing capabilities to ML training frameworks. It shares the same limitations as Spark.
MillWheel [2] is a stream processing system that offers efficient reconfiguration and failover by combining decentralized physical logging with a centralized load-balancer off the critical path. It offers sophisticated APIs for real-time processing, including timers, watermarks, etc. In contrast, Ray Data targets offline processing, uses lineage-based recovery to avoid data logging, and the centralized scheduler dispatches all tasks for a global view and finer control over resources.
Other systems have explored building distributed data processing frameworks on top of task-parallel systems such as Ray [27, 44], CIEL [29] and Dask [37]. However, all of these systems are CPU-centric and do not consider memory or heterogeneous compute.
@@
The scheduling problem described in Section 4.3 is most similar to the generalized processor sharing [33] problem. Our solution is inspired by the weighted fair queueing algorithm [6, 12]. The differences are (1) the flows in network scheduling are independent of each other, whereas operators in a data pipeline have dependencies, (2) multiple resource types, and (3) the packet processing time is usually fixed, whereas data operator processing times are unpredictable. (1) is important in that operators may produce significant intermediate data, which may not be released until the downstream operator executes.
Recent scheduling works attempt to adapt fair queuing and autotuning to heterogeneous resource environments. Dominant resource fair queuing [15] addresses the problem of (2) but not (1) or (3). tf.data [30] introduces an autotuning algorithm that uses gradient descent to find the parallelism for each operator that reduces end-to-end latency. We instead aim to maximize overall throughput; the partition size may be used to adjust end-to-end latency.
Streaming data loaders for ML.
Both PyTorch [34] and TensorFlow (tf.data [30]) provide data loaders optimized for map-style transforms for ML training. tf.data uses multithreading while PyTorch DataLoader uses multiprocessing. tf.data automatically and deterministically shards the dataset, while PyTorch DataLoader requires the user to shard the dataset themselves. However, both are single-node systems colocated with a GPU trainer and share similar limitations: they cannot execute multi-node, dataset sharding must be done before execution, and the training job fate-shares with the data loader.
Cloud-native streaming data loaders include MosaicML Streaming [21] and DeepLake [18]. The common innovations in these libraries are a specialized data format optimized for tensor storage and querying, and streaming data loading directly from cloud storage. They use Python multiprocessing for execution, which can be swapped out for Ray Data. This would decouple data loading from ML training/inference processes, thus allowing more flexible control of CPU parallelism and memory buffers and scale-out to heterogeneous nodes.
@@ 7 Discussion
One benefit of building Ray Data as a Ray library rather than a monolithic system such as Spark or Flink is that since the Ray core is in C++, Ray Data can support many frontend languages, including the lingua franca for ML data loading and transformations, Python. In contrast, systems built on non-native languages show high overheads when a Python frontend is used (e.g., Flink in § 5.1.1). Second, modifying the Ray Data scheduler is convenient, as it is written in the frontend and does not require re-compiling Ray [25].
Ray Data enables another key opportunity in future data processing systems: dynamic query planning. In this work, we present an online scheduler, but still make certain planning decisions statically, including the number of input partitions (§ 4.1), and the user specifies the initial cluster shape. We envision a fully autotuning and autoscaling system that can cohesively re-plan the application and resize the cluster.
@@ Conclusion.
With the rise of large language models, even modalities such as text that are traditionally not compute-intensive to load and process may now require expensive deduplication [31], GPU-based embedding computation, and joins with image or video data (Figure 1(b)). In addition, as the parallelism strategies used in inference and training pipelines become more complex, future data processing systems must also support more flexible APIs for sharding and sharing data.
In general, we believe that ML systems will continue to grow in the complexity of their data processing needs, as evidenced by trends such as test-time training [39, 14], retrieval-augmented generation [24], and multimodal models [40, 32]. To keep up with this demand, we must build more flexible, heterogeneity-aware, and scalable data processing systems.
@@ A.1 Fractional Parallelism
In this microbenchmark, we demonstrate that the streaming batch model can maximize the resource utilization when fractional parallelism is required. Consider a two-stage data pipeline, in which the first stage takes 1 second on average, and the second stage takes 2 seconds. Ideally, the operator parallelisms should be set as $2:1$ to balance the throughput. In traditional stream processing systems such as Flink, this is unattainable on a 8-CPU machine, because it requires setting the operator parallelisms to be 2.67 and 5.33, respectively. Since these systems allocate executors to operators statically, they cannot support fractional parallelism. In contrast, the streaming batch execution model allows Ray Data to multiplex executors for both stages dynamically during run time. The Ray Data scheduler can dynamically start a task for either stage in order to balance the throughput, effectively achieving a parallelism ratio of $2:1$ over time. Figure 11 shows that when comparing to a static allocation of 4–4 executors for each stage, the dynamic allocation increases the utilization of the execution slots, manifested as fewer bubbles in the schedule, and 19% faster job completion time.
@@ B.1 Algorithm
The solver starts from the initial state (time 0, all executors idle, buffer empty, and all tasks pending). For each state, it generates the next state by emulating the execution: advancing the tick, updating executor states, updating the progress of running tasks, updating the memory buffer, etc. The number of next states is determined by the size of the set of all possible scheduling actions, which is the power set of all possible scheduling primitives. A scheduling primitive would be “schedule the next task operator $i$ onto executor $j$.”
@@
For the scheduling microbenchmark in Section 5.3.1, the solver finds the optimal schedule with a total run time of 153 seconds.

### v5

@@ 5.3.1 Memory-aware pipelining
For heterogeneous applications under memory pressure, batch processing systems are unstable. Ray Data is as stable as stream processing systems, due to its dynamic repartition, and it can further leverage run-time profiling.
@@ 5.3.2 Overhead of partitioning
Ray Data uses dynamic repartitioning with a centralized scheduler. Figure˜10(a) evaluates Ray Data’s system overhead by measuring throughput vs. partition size on a 2-stage synthetic pipeline. We use $8192\times$1 MB input rows, and simulate 10 ms processing time per row per stage. The smallest partition sizes incur overhead from RPCs and bookkeeping, and the largest result in poor load-balancing. To strike a balance, Ray Data’s default target partition size is 128 MB.
Figure 10: (a) Effect of partition sizes on throughput in Ray Data. (b) Scalability experiment showing throughput over number of nodes compared to using raw Ray tasks or generator tasks (Ray Generator).
5.3.3 Scalability
To evaluate the scalability of Ray Data, we run an empty workload which creates then consumes a large dataset on the cluster. We use 1 m8i.4xlarge (16 vCPU, 64GiB RAM) as the head node and up to 32 m8i.2xlarge (8 vCPU, 32GiB RAM) worker nodes. The size of the dataset is proportional to the number of nodes in the cluster, with 5GB per node. We compare Ray Data against two baselines: Ray and Ray Generator. Ray Data uses the default partition size of 128MB. Ray spawns tasks that each return 128MB of data, then immediately passes each result to a consumer task. Ray Generator does the same, except that it uses the generator tasks described in Section˜4.2.1.
[…]
@@ 6 Related Work
Other systems that aim to unify the batch and stream models include Naiad [27], Flink’s batch execution mode [7], and Spark’s Discretized Streams (D-Streams) [44]. Naiad shows that the stream processing model is suitable for producing results both incrementally and in bulk, but it does not support dynamic resource assignment. Flink’s batch mode executes using synchronous stages, with the same limitations as other batch processing systems (Section˜2.2).
Spark D-Streams and Drizzle [40] are batch systems that execute over infinite data streams. They partition the input stream into “microbatches”, each executed as a distinct job and typically one at a time. Internally, each microbatch job uses synchronous stage execution. If different microbatches were pipelined, it can be viewed as a form of the streaming batch model, but with three limitations: (1) the pipeline granularity is a microbatch job instead of a partition within a microbatch, (2) each microbatch is treated as a separate job, so scheduling and data movement across microbatches is challenging, and (3) data partitioning is still static.
MillWheel [1] is a stream processing system that offers efficient reconfiguration and failover by combining decentralized physical logging with a centralized asynchronous load-balancer. It offers sophisticated APIs for real-time processing, including timers, watermarks, etc. In contrast, Ray Data targets offline processing, uses lineage-based recovery to avoid data logging, and the centralized scheduler dispatches all tasks for a global view and more control over resources.
Other efforts [3, 2] aim to unify streaming and batch execution but only at the API layer.
@@
The scheduling problem described in Section˜4.3 is most similar to the generalized processor sharing [31] problem. Our solution is inspired by the weighted fair queueing algorithm [6, 9]. The differences are (1) the flows in network scheduling are independent of each other, whereas operators in a data pipeline have dependencies, (2) we consider multiple resource types, and (3) the packet processing time is usually fixed, whereas data operator processing times are unpredictable.
Recent scheduling works attempt to adapt fair queuing and autotuning to heterogeneous resource environments. Dominant resource fair queuing [13] addresses the problem of (2) but not (1) or (3). tf.data [28] introduces an autotuning algorithm that uses gradient descent to find the parallelism for each operator that reduces end-to-end latency.
Data loaders for ML training.
Both PyTorch [32] and TensorFlow (tf.data [28]) provide map-style data loaders for ML training. tf.data automatically shards the dataset, while PyTorch DL requires the user to shard the dataset themselves. Both are single-node systems colocated with a GPU trainer and share similar limitations: they cannot leverage CPU-only nodes, load-balancing must be done by sharding before execution, and the training job fate-shares with the data loader.
tf.data service [5] improves upon these by enabling disaggregated CPU-based preprocessing for training. Ray Data provides a similar set of features and additionally provides exactly-once semantics (as opposed to at-most-once), GPU-based preprocessing, and batch inference support. Several recent works propose query optimizers for training preprocessing [15, 45]. These are complementary to Ray Data.
@@ 7 Discussion
Achieving native performance with a Python system is an important practical consideration for the modern ML ecosystem. One benefit of building Ray Data as a Ray library rather than as a monolithic system such as Spark or Flink is the ability to write Ray Data in pure Python. The Ray core is in C++, which avoids some of the overheads of exposing a Python frontend compared to systems built on non-native languages. Another benefit is that it is easy to modify the Ray Data scheduler, as it is written at the library level.
Ray Data enables another key opportunity in future data processing systems: dynamic query planning. In this work, we present an online scheduler, but still make certain planning decisions statically, including the number of input partitions (Section˜4.1), and the user specifies the initial cluster shape. In the future, we envision a fully autotuning and autoscaling system that can jointly configure the application and cluster.
@@ Conclusion.
With the rise of LLMs, even modalities such as text may now require expensive deduplication [29], retrieval (Section˜5.1.1), GPU-based encoding, and joins with image or video data (Figure˜1(b)). Also, as the parallelism strategies used in inference and training pipelines become more complex, future data processing systems must support more flexible APIs for sharding and reducing data copies.
We believe that ML systems will continue to grow in the complexity of their data processing needs, as evidenced by trends such as test-time training [36, 12], RAG [23], and multimodal models [37, 30]. To keep up with this demand, we must build more flexible, heterogeneity-aware, and scalable data processing systems.
@@ A.1 Fractional Parallelism
In this microbenchmark, we demonstrate that the streaming batch model can maximize the resource utilization when fractional parallelism is required. Consider a two-stage data pipeline, in which the first stage takes 1 second on average, and the second stage takes 2 seconds. Ideally, the operator parallelisms should be set as $2:1$ to balance the throughput. In traditional stream processing systems such as Flink, this is unattainable on a 8-CPU machine, because it requires setting the operator parallelisms to be 2.67 and 5.33, respectively. Since these systems allocate executors to operators statically, they cannot support fractional parallelism. In contrast, the streaming batch execution model allows Ray Data to multiplex executors for both stages dynamically during run time. The Ray Data scheduler can dynamically start a task for either stage in order to balance the throughput, effectively achieving a parallelism ratio of $2:1$ over time. Figure˜11 shows that when comparing to a static allocation of 4–4 executors for each stage, the dynamic allocation increases the utilization of the execution slots, manifested as fewer bubbles in the schedule, and 19% faster job completion time.
@@ B.1 Algorithm
The solver starts from the initial state (time 0, all executors idle, buffer empty, and all tasks pending). For each state, it generates the next state by emulating the execution: advancing the tick, updating executor states, updating the progress of running tasks, updating the memory buffer, etc. The number of next states is determined by the size of the set of all possible scheduling actions, which is the power set of all possible scheduling primitives. A scheduling primitive would be “schedule the next task operator $i$ onto executor jj.”
@@
For the scheduling microbenchmark in Section˜5.3.1, the solver finds the optimal schedule with a total run time of 153 seconds.